# ENSICAEN, CNRS, GREYC

### 06 Mar 2024 v5

@@ I Introduction
When analyzing the relationship between feature vectors or concurrent signals, correlation and covariance matrices are a useful tool. Such matrices are at least Positive Semi-Definite, and often fully Symmetric Positive Definite (SPD). The set of $n\times n$ SPD matrices ($SPD(n)$) is a non-Euclidean, Riemannian (i.e. metric) manifold, and the regular Euclidean operations of most Neural Network (NN)-based models seldom preserve that geometric structure, introducing deformations such as the “swelling effect” [1]. Structure-preserving NN-based approaches have been introduced [2, 3], deriving their layers from one of two geodesic-defining metrics on $SPD(n)$. Affine invariant metrics offer the best properties, but present computational challenges (e.g. no closed-form formula for averaging) [4]. LogEuclidean metrics are less isotropic, but still prevent swelling while being easier to compute [1]. With $A,B\in SPD(n)$, we chose this LogEuclidean distance:
@@
In this paper, we present a structure-preserving self-attention mechanism applicable to sequences of SPD matrices, derived from the aforementioned LogEuclidean metric. We embed said mechanism into a Transformer-based architecture, and apply it to a biomedical classification problem. Transformer-based technology has exploded in popularity ever since its introduction in [5], with self-attention mechanisms being applied to very different problems. With regards to Riemannian geometry, innovations seem centered around the computation and application of attention maps, specifically. For instance, Konstantinidis et al. [6] combine the standard attention maps with Grassmann and SPD manifold-valued maps, to enrich their computer vision model’s descriptive capabilities. By contrast, both He et al. [7] and Li et al. [8] developed architectures to analyze 2D-manifold-valued data in 3D space, the former achieving rotational equivariance with respect to surfaces on the manifold and the latter developing two geodesic distances applicable to point clouds, and building attention maps from these distances. More generally, Kratsios et al. [9] provide a mathematical framework to apply attention mechanisms on a variety of constrained sets, including manifolds. While the latter approaches share our interest in preserving geometric information, little to no focus is given to a Transformer’s other components. As far as we are aware, ours is the only approach to apply structure-preserving Transformers to SPD manifold-valued data.
@@ II-A Structure-Preserving Multihead Attention (SP-MHA)
with $\mathcal{L}_{Q}(\cdot)$ and $\mathcal{L}_{K}(\cdot)$ banks of $h$ linear maps from $\mathbb{R}^{d(m)}$ to $\mathbb{R}^{\frac{d(m)}{h}}$, $sm(\cdot)$ the softmax function, and $C(\cdot)$ the weighted linear combination of the $h$ post-softmax attention maps. Here, the computation of attention maps (small-dashed black rectangle in Figure 1) remains identical to L-MHA. However, their application to V (large-dashed red rectangle in the figure) only requires a matrix multiplication, i.e. linear combinations of V’s tokens weighted by the combined attention map. As such, the SP-MHA bloc does not threaten our tokens’ vector space geometry.
@@ II-B Triangular linear maps
III Application to EEG Sleep Staging
III-A Improving stage-wise classification
@@
Seraphim et al. [15] hypothesized that an analysis through functional connectivity - the activation correlations between different brain regions [23] - enhances stage-wise performance. Such an analysis was first done by Jia et al. [13], using epoch-wise graph learning to estimate said connectivity and sequence-wise spatio-temporal graph NNs to compare them. By contrast, Seraphim et al. estimate it through covariance matrices. Their two-step model uses standard Transformer encoders at each step, reminiscent of [22]. Each input epoch is described as a multichannel timeseries of SPD matrices, which are then tokenized bijectively. However, their approach does not guarantee the preservation of their data’s SPD structure, as they operate a channel-wise concatenation of their tokens, in addition to the concatenations found within their encoders (Section II-A). Hence, we propose a Transformer-based model capable of analyzing EEG-derived functional connectivity through SPD matrices without sacrificing the SPD structure of our data throughout the analysis.

### 08 Mar 2024 v6

@@ I Introduction
When analyzing the relationship between concurrent signals, covariance matrices are a useful tool, with applications in fields like Brain-Computer Interfaces (BCI) [1] and evolutionary computation [2]. By construction, they are rich in information, illustrating the relationship between signals while still encoding for signal-wise information on their diagonal. Such matrices are at least Positive Semi-Definite, and often fully Symmetric Positive Definite (SPD). The set of $n\times n$ SPD matrices ($SPD(n)$) is a non-Euclidean, Riemannian (i.e. metric) manifold, and the regular Euclidean operations of most Neural Network (NN)-based models seldom preserve that geometric structure, introducing deformations such as the “swelling effect” [3]. Structure-preserving NN-based approaches have been introduced [4, 5], deriving their layers from one of two geodesic-defining metrics on $SPD(n)$. Affine invariant metrics offer the best properties, but present computational challenges (e.g. no closed-form formula for averaging) [6]. LogEuclidean metrics are less isotropic, but still prevent swelling while being easier to compute [3]. With $A,B\in SPD(n)$, we chose this LogEuclidean distance:
@@
In this paper, we present a structure-preserving self-attention mechanism applicable to sequences of SPD matrices, derived from the aforementioned LogEuclidean metric. We embed said mechanism into a Transformer-based architecture, and apply it to a biomedical classification problem. Transformer-based technology has exploded in popularity ever since its introduction in [7], with self-attention mechanisms being applied to very different problems. With regards to Riemannian geometry, innovations seem centered around the computation and application of attention maps, specifically. For instance, Konstantinidis et al. [8] combine the standard attention maps with Grassmann and SPD manifold-valued maps, to enrich their computer vision model’s descriptive capabilities. By contrast, both He et al. [9] and Li et al. [10] developed architectures to analyze 2D-manifold-valued data in 3D space, the former achieving rotational equivariance with respect to surfaces on the manifold and the latter developing two geodesic distances applicable to point clouds, and building attention maps from these distances. More generally, Kratsios et al. [11] provide a mathematical framework to apply attention mechanisms on a variety of constrained sets, including manifolds. While the latter approaches share our interest in preserving geometric information, little to no focus is given to a Transformer’s other components. To the best of our knowledge, our approach is the only one applying structure-preserving Transformers to SPD manifold-valued data.
@@ II-A Structure-Preserving Multihead Attention (SP-MHA)
with $\mathcal{L}_{Q}(\cdot)$ and $\mathcal{L}_{K}(\cdot)$ banks of $h$ linear maps from $\mathbb{R}^{d(m)}$ to $\mathbb{R}^{\frac{d(m)}{h}}$, $sm(\cdot)$ the softmax function, and $C(\cdot)$ the weighted linear combination of the $h$ post-softmax attention maps. Here, the computation of attention maps (small-dashed black rectangle in Figure 1) remains identical to L-MHA. However, their application to V (large-dashed red rectangle in the figure) only requires a matrix multiplication, i.e. linear combinations of V’s tokens weighted by the combined attention map. As such, the SP-MHA bloc does not compromise our tokens’ vector space geometry.
@@ II-B Triangular linear maps
III Application to EEG Sleep Staging
III-A Improving stage-wise classification
@@
Beyond sleep staging, EEG signals are also utilized in BCI (Section I), where they are often analyzed through the lens of functional connectivity - the activation correlations between different brain regions [25]. Automatic sleep staging through functional connectivity was first investigated by Jia et al. [23], using epoch-wise graph learning to estimate said connectivity and sequence-wise spatio-temporal graph NNs to compare them. By contrast, Seraphim et al. [22] estimate it through covariance matrices, as is commonly done in BCI [1]. Their two-step model uses standard Transformer encoders at each step, reminiscent of [21]. Each input epoch is described as a multichannel timeseries of SPD matrices, which are then tokenized bijectively. However, their approach does not guarantee the preservation of their data’s SPD structure, as they operate a channel-wise concatenation of their tokens, in addition to the concatenations found within their encoders (Section II-A). Hence, we propose a Transformer-based model capable of analyzing EEG-derived functional connectivity through SPD matrices without sacrificing the SPD structure of our data throughout the analysis.